ariana123:

A white-tailed deer can sprint at speeds up to 30 miles per hour. American bison can run at speeds up to 3,520 feet per minute. Which animal is faster and by how many miles per hour? There are 5,280 feet in one mile.

justjm:

Convert the values if necessary to comparable measures. You can convert the 3520 ft/minute to miles/hour. Use dimensional analysis in doing so.

ariana123:

alright , thank you !

justjm:

np, to get you started \(\frac{3520~ft}{1~minute} * \frac{60~minutes}{1~hour} * \frac{1~mile}{5280~ft} \) This will get the 3520 ft/minute to mph. Note how the units cancel